# The SDSC Satellite Reverse Proxy Service for Launching Secure Jupyter Notebooks on High-Performance Computing Systems

### v1

@@ I Introduction
The Jupyter Project provides several tools and services including the Notebook Server (which launches Jupyter Notebooks, and JupyterLabs) and JupyterHub (which spawns, manages, and proxies multiple notebooks for multiple users) [1]. Because of their popularity, these tools and services are deployed on a large variety of High-Performance Computing Systems (HPC) resources [2, 3, 4, 5]. SDSC continues to support hosting Jupyter services on its HPC systems. However, as part of investigating the best methods for hosting JupyterHub and Jupyter Notebook services on SDSC HPC systems, we discovered several challenges and concerns in the area of security and resource usage issues.
@@ III Using Satellite
For more details on using Satellite on the SDSC Expanse cluster, see the Notebooks 101 User Guide [6], and for example notebooks that run on SDSC HPC systems, for both CPU and GPU devices, see the Notebook-Examples GitHub repository [7].
@@ IV Conclusions and Future Plans
Future plans also include the addition of unit testing using Travis CI Test system (a service used to build and test software projects hosted on GitHub and Bitbucket [8]) ; exploring the use of Internal public-key-infrastructure for enabling TLS/HTTPS between Satellite and proxied applications; the development of usage metrics; and updating the Satellite service installation and deployment architecture.

### v2

@@ I Introduction
The Jupyter Project provides several tools and services including the Notebook Server (which launches Jupyter Notebooks, and JupyterLabs) and JupyterHub (which spawns, manages, and proxies multiple notebooks for multiple users) [ProjectJupyterHome]. Because of their popularity, these tools and services are deployed on a large variety of High-Performance Computing Systems (HPC) resources [Milligan-PEARC17, Zonca-PEARC18, Nicklas2018, Stubbs2020]. SDSC continues to support hosting Jupyter services on its HPC systems. However, as part of investigating the best methods for hosting JupyterHub and Jupyter Notebook services on SDSC HPC systems, we discovered several challenges and concerns in the area of security and resource usage issues.
@@ III Using Satellite
For more details on using Satellite on the SDSC Expanse cluster, see the Notebooks 101 User Guide [Notebooks101], and for example notebooks that run on SDSC HPC systems, for both CPU and GPU devices, see the Notebook-Examples GitHub repository [github-ntbk-examp].
@@ IV Conclusions and Future Plans
Future plans also include the addition of unit testing using Travis CI Test system (a service used to build and test software projects hosted on GitHub and Bitbucket [TravisCI]) ; exploring the use of Internal public-key-infrastructure for enabling TLS/HTTPS between Satellite and proxied applications; the development of usage metrics; and updating the Satellite service installation and deployment architecture.